OpenStudy (anonymous):

find the distance between points M(6,16) and Z(1,14) to the nearest tenth

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ M&({\color{red}{ 6}}\quad ,&{\color{blue}{ 16}})\quad Z&({\color{red}{ 1}}\quad ,&{\color{blue}{ 14}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

so would it be\[\sqrt{29}\]